# Page Pages

### v1

@@ 8 Summary and Conclusions
In this work, we studied correlation functions along three axes: model uncertainty, differentiability, and surrogates. To better understand the model uncertainty caused by clustering approximations, we used $\rho$ statistics as a case study. We made our assignment of objects to clusters probabilistic, and computed the mean and standard deviation of the resulting correlation estimates. Our analysis compared the uncertainty estimate to traditional bootstrapping methods. Our next experiment adapted our main algorithm to be automatically differentiable and applied this to the setting of gravitational simulations. Three different approaches of doing this were proposed, and of those, approach $c$ proved to be the most stable. Still, the calculation of the gradient was slow and approximate, motivating the use of surrogate methods. To that end, we ended by exploring surrogate solutions. We extended the work of Pandya et al. (2024, 2025), showing how we can exploit the gradients of our surrogate model to optimize over astrophysical model parameters. Specifically, we studied galaxy intrinsic alignment, and found which astrophysical model parameters were most responsible for high / low correlations (Figures 12-13).
@@ acknowledgements
E.B. thanks Quinn Arbolante, Jacob Ginesin, Luisa Li, Eric Zelikman for helpful discussions. E.B. also thanks Gregory Shomo for helpful guidance with the discovery computing cluster. S.P. acknowledges support from the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, https://iaifi.org). Support for the COSMOS-Web survey was provided by NASA through grant JWST-GO-01727 and HST-AR-15802 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 526555. This work was made possible by utilizing the CANDIDE cluster at the Institut d’Astrophysique de Paris, which was funded through grants from the PNCG, CNES, DIM-ACAV, and the Cosmic Dawn Center and maintained by Stephane Rouberol. Further support was provided by Research Computing at Northeastern University.
@@ Appendix C Jacobians
Figure 18: Jacobians associated with each of the $7$ IA parameters and $20$ output bins from IA-Emu.
Glossary

### v3

@@ 8 Summary and Conclusions
In this work, we studied correlation functions along three axes: model uncertainty, differentiable forward models, and differentiable surrogates. To better understand the model uncertainty caused by clustering approximations, we used $\rho$ statistics as a case study. We made our assignment of objects to clusters probabilistic, and computed the mean and standard deviation of the resulting correlation estimates. Our analysis compared the uncertainty estimate to traditional bootstrapping methods. Our next experiment adapted our main algorithm to be automatically differentiable and applied this to the setting of gravitational simulations. Three different approaches of doing this were proposed, and of those, approach $c$ proved to be the most stable. Still, the calculation of the gradient was slow and approximate, motivating the use of surrogate methods. To that end, we ended by exploring surrogate solutions. We extended the work of Pandya et al. (2024, 2025), showing how we can exploit the gradients of our surrogate model to optimize over astrophysical model parameters. Specifically, we studied galaxy intrinsic alignment, and found which astrophysical model parameters were most responsible for high / low correlations (Figures 12-13).
@@ acknowledgements
E.B. thanks Quinn Arbolante, Jacob Ginesin, Luisa Li, Eric Zelikman for helpful discussions. E.B. is supported by a College of Science Dean’s Office Research grant. E.B. also thanks Gregory Shomo for helpful guidance with the discovery computing cluster. S.P. acknowledges support from the National Science Foundation under Cooperative Agreement PHY-2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, https://iaifi.org). Support for the COSMOS-Web survey was provided by NASA through grant JWST-GO-01727 and HST-AR-15802 awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 526555. This work was made possible by utilizing the CANDIDE cluster at the Institut d’Astrophysique de Paris, which was funded through grants from the PNCG, CNES, DIM-ACAV, and the Cosmic Dawn Center and maintained by Stephane Rouberol. Further support was provided by Research Computing at Northeastern University.
@@ Appendix C Jacobians
Figure 18: Jacobians associated with each of the $7$ IA parameters and 2020 output bins from IA-Emu.